# MobCLIP: Learning General-purpose Geospatial Representation at Scale

### v3

@@ 5 Conclusion
In this work, we present MobCLIP, a general-purpose framework for geospatial representation learning at scale. Centered around a billion-scale mobility graph and enriched with multimodal alignment—including POIs, remote sensing imagery, and tabular data—MobCLIP learns unified region embeddings that generalize across diverse downstream tasks. Extensive experiments on 11 tasks spanning social, economic, and natural domains demonstrate that MobCLIP consistently outperforms existing state-of-the-art methods. Moreover, our findings echo the scaling laws observed in other domains: as the spatial coverage of training data increases from local to national scale, representation quality and task performance improve substantially.
Despite strong empirical results, limitations remain. The scale issue in spatial representation learning is not fully addressed. We conducted a sensitivity analysis to show that MobCLIP is robust over different spatial scales (Appendix A.3). Second, high-quality human mobility data may be difficult to obtain in some regions for obvious privacy concerns. To resolve this issue, we generate and open-source a distilled version of MobCLIP that is free from privacy concerns. Specifically, we train a multi-layer perceptron (MLP) to map continuous geographic coordinates into high-dimensional embeddings, supervised by MobCLIP’s grid-level outputs. This approach not only enables users to retrieve high-quality region embeddings for any arbitrary coordinate, free from privacy concerns, but also removes the dependency on H3-based lookup. Detailed implementation of the distillation approach is provided in Appendix A.1.
@@ A.1 Detailed Implementation of the Distillation
For model distillation, we fit a surrogate Multi-Layer Perceptron (MLP) network to estimate pretrained region embeddings from MobCLIP. For each H3 cell, the input for the MLP is a 1024-d vector, which is a SIREN encoding [33] of the centroid coordinate of the cell. The output for the MLP is a 128-d vector from MobCLIP. The MLP is configured with 8 hidden layers, with a dimension of 512 each. The model is trained with an MSE loss function to minimize the deviation of estimated embeddings from our true embeddings. The learning rate is set to 0.005 for 5000 epochs. Final training loss is 0.0857. The downstream performance of the distilled model is shown in Table 3. Note that the performance of the distilled model (MobCLIPd) tends to improve as the number of MLP layers and the size of the hidden dimensions increase. We here demonstrate the distillation effectiveness preliminarily using the model size mentioned above.
Table 3: Performance comparison between MobCLIPd and baselines.
Social Tasks Economic Tasks Natural Tasks
POP EDU ELD HSR CRI NTL HOU ENE COS ELE FOR
SatCLIP $0.523$ $0.633$ $0.676$ $0.735$ $0.391$ $0.334$ $0.655$ $-0.069$ $0.443$ $0.963$ $0.806$
MobCLIP $0.829$ $0.853$ $0.812$ $0.810$ $0.759$ $0.615$ $0.703$ $0.719$ $0.909$ $0.900$ $0.797$
MobCLIPd $0.730$ $0.835$ $0.804$ $0.816$ $0.580$ $0.517$ $0.693$ $0.742$ $0.729$ $0.906$ $0.785$
A.2 Results Table for the Scaling Behavior
@@
A.3 Sensitivity to H3 Grid Resolution
We extend our study at H3 resolution level 6 to its adjacent levels, 5 and 7, to evaluate the model’s performance across different spatial scales. Compared to level 6 H3 cells, which cover an average area of 36.13 km², cells at levels 5 and 7 cover approximately 253.61 km² and 5.18 km², respectively. The East China region is selected for testing to facilitate ease of operation without losing generalizability. We reconstruct the training dataset for levels 5 and 7 from scratch, including forming trajectory-based mobility graphs, gathering satellite imagery and POIs for encodings, and processing demographic distributions data. Three separate experiments are conducted, with corresponding sets of downstream data processed for evaluation. Since cities in East China are limited in quantity, we choose Ridge Regression for electricity prediction across all levels to ensure consistency. For the remaining tasks, LightGBM is adopted. Table 5 demonstrates the model’s performance with varying resolutions. For grid-level predictions, the performance tends to decline as the level of granularity increases. This is because, at finer resolutions, the data becomes more heterogeneous and exhibits greater variations. However, an exception is observed with population prediction, which follows a reversed pattern. This might be explained by the fact that population is a localized variable that demonstrates stronger patterns at smaller spatial scales. On the other hand, aggregating spatial units smooths out local fluctuations and emphasizes important features in the embedding, as evidenced by the improvement in the model’s performance on county- and city-level tasks with higher resolution levels.
Table 5: Sensitivity of model performance using different grid resolutions.

### v4

@@ 5 Conclusion and Discussion
This work introduces MoRA, a mobility-as-the-backbone framework for geospatial representation learning. Centered around a billion-scale mobility graph and enriched with multimodal alignment (including POIs, remote sensing imagery, and tabular data), MoRA learns unified region embeddings that generalize across diverse downstream tasks. Extensive experiments on 9 socio-economic tasks demonstrate MoRA’s consistent outperformance of state-of-the-art methods. Moreover, our findings echo the scaling laws observed in other domains: expanding training data’s spatial coverage from local to national scales significantly enhances both representation quality and task performance.
Despite strong empirical results, limitations remain. The scale issue in spatial representation learning is not fully resolved. Nevertheless, our sensitivity analysis indicates that MoRA is robust across spatial scales (Appendix A.6). Second, although MoRA itself as a framework is dataset-agnostic and can be adapted to alternative, publicly available movement data sources (e.g., Internet data or mobile phone data), high-quality human mobility data may be difficult to obtain in some regions. In such cases, we suggest a neighborhood-based variant of our model as a viable and effective alternative that produces high-quality geospatial representation. Both the code and the distilled model will be released at: https://github.com/ylzhouchris/MoRA.
6 Ethics statement
This work adheres to the ICLR Code of Ethics. Recognizing the privacy considerations associated with human mobility data, we rigorously minimize associated risks through a four-stage anonymization-aggregation-learning-distillation pipeline:
  1. 1.
    Anonymization: The mobility data is derived from the locations of offline stores on the WeChat Pay platform. Grids are linked when stores in different regions are visited by the same user within a week, and these links are aggregated into an inter-grid network. No individual-level information is retained.
  2. 2.
    Aggregation: Raw mobility data is aggregated at a much coarser grid level, eliminating any individual-level information.
  3. 3.
    Representation Learning: Embeddings inherently generalize patterns.
  4. 4.
    Model Distillation: We share a distilled model, ensuring that reconstructing the original data or tracing back to specific individuals is practically impossible.
This aligns with established privacy-preserving practices in mobility research. The disclosed embeddings contain no personally identifiable information, and our distillation adds an additional layer of critical protection.
7 Reproducibility statement
We have made every effort to ensure a high level of reproducibility of our work. Specifically, we:
  • Release the complete codebase at https://github.com/ylzhouchris/MoRA, including data preprocessing pipelines, model architecture, and code for downstream task evaluation.
  • Describe the training implementation in detail in Section 4.1, including training steps, hardware details, model hyperparameters and so on.
  • Make the pretrained model weights available by open-souring a high-performance distilled version of our model, enabling researchers to immediately benefit from MoRA’s capabilities via simple, direct queries.
  • Establish and release benchmark downstream dataset, accompanied by detailed descriptions of task properties and visualizations of their geographical distributions in Appendix A.2.
@@ A.1 Detailed Implementation of the Distillation
For model distillation, we fit a surrogate Multi-Layer Perceptron (MLP) network to estimate pretrained region embeddings from MoRA. For each H3 cell, the input for the MLP is a 1024-d vector, which is a SIREN encoding (Rußwurm et al., 2024) of the centroid coordinate of the cell. The output for the MLP is a 128-d vector from MoRA. The MLP is configured with 8 hidden layers, with a dimension of 512 each. The model is trained with an MSE loss function to minimize the deviation of estimated embeddings from our true embeddings. The learning rate is set to 0.005 for 5000 epochs. Final training loss is 0.0857. The downstream performance of the distilled model is shown in Table 4. Note that the performance of the distilled model (MoRAd) tends to improve as the number of MLP layers and the size of the hidden dimensions increase. We here demonstrate the distillation effectiveness preliminarily using the model size mentioned above.
Table 4: Performance comparison between MoRAd and baselines.
Social Tasks Economic Tasks
POP EDU ELD HSR CRI NTL HOU ENE COS
SatCLIP 0.523 0.633 0.676 0.735 0.391 0.334 0.655 -0.069 0.443
MoRA 0.829 0.853 0.812 0.810 0.759 0.615 0.703 0.719 0.909
MoRAd 0.730 0.835 0.804 0.816 0.580 0.517 0.693 0.742 0.729
A.2 Details on Downstream Datasets and Spatial Distribution
The data sources summarized in Table 1 are primarily drawn from WorldPop (Bondarenko et al., 2020), MOSAIKS (Rolf et al., 2021), and the 2020 China Census Year Book from the National Bureau of Statistics of China (NBS, 2020). Additionally, ENE (energy consumption) is compiled from various local statistical yearbooks. The total energy consumption of one city/prefecture is computed by converting different forms of energy usage into tonnes of standard coal equivalent. We also include an unavailable-to-the-public dataset, COS (offline consumption amount) from Tencent WeChat Pay, to evaluate whether the embeddings can predict regional consumption indices. For the crime dataset, using the original data from (Zhang et al., 2025), we aggregate cases over a five-year interval, add 1 and apply a log transform to the labels.
Figure 4 depicts the geographical distribution of the downstream task datasets. Overall, samples are widely spread across China. Note, however, that the west contains vast plateaus and deserts (e.g., the Tibetan Plateau) with minimal human activity, leading to sparse coverage for several tasks. In regions with active human presence, samples are distributed fairly evenly. Importantly, these maps reflect only the task-specific data availability—some sparsity stems from the nature of the tasks or their collection processes—and is independent of our region embeddings, which are uniformly distributed across populated areas of China.
Figure 4: The geographical distribution of various downstream dataset samples.
A.3 Baselines
We benchmark our framework against two types of baselines: (i) feature-based extractors that enable direct, coordinate-based, training-free embedding queries through pretraining or analytical transformation functions, and (ii) methods that require either retraining or the preparation of a strictly matching input dataset. Since the first category aligns closerly with our approach and is covered in details in Related Works, the discussion here focuses on the second category. A notable line of work in this category adopts graph-based multi-view representation learning: for each view (e.g., mobility, POI), a graph is constructed based on attribute similarity, and the resulting node embeddings are fused (e.g., via attention mechanisms) to produce unified region representations. Contrastive-based methods—mostly formulated between two views or modalities of the data—are also widely used for city-scale region representation learning. Besides, there are early attempts for building region representation fundation model through multi-city pretraining or distillation from LLMs and VLMs. Accordingly, we select three representative state-of-the-art works from each category as baselines:
  • HREP constructs a heterogeneous graph from three types of regional data, namely, human mobility, POIs and geographic neighboring relationship, and learns region embeddings using a relation-aware GCN and attention-based cross-view fusion.
  • ReCP leverages views from POI and human mobility data, applies contrastive learning at both intra-view and inter-view levels to enhance the cross-view consistency, and finally fuse the two views via concatenation.
  • ReFound integrates multi-source grospatial data through a transformer-based structure and distills knowledge from language, visual foundation models joinlt to enhance its generalization capabilities. The model is pretrained in five major cities in China to acquire general geospsatial knowledge.
The baselines for HREP and ReCP require re-training, whereas ReFound provides pre-trained weights for inference. However, because the datasets for both input features and downstream evaluation are unavailable for all three models, we adapt their codebases and released weights to our datasets in the same study area. Specifically, we find and align required data with each model’s input format, re-train HREP and ReCP, and run inference with ReFound.
A.4 Results Table for the Scaling Behavior
Table 5 reports task-wise performance for the scaling experiments described in Section 4.3 , which also serves as the source data for Figure 2.
@@ A.4 Results Table for the Scaling Behavior
A.5 Ablation on modalities
We ablate each of the three auxiliary modalities, i.e., POIs, satellite imagery, and demographics, to assess their contributions to the model performance on human-centric inference. As shown in Table LABEL:table:ablation_on_modalities_on_modalities, the removal of deomographics information consistenly results in the largest performance decrease across nearly all tasks, demonstrating its unique advantages in capturinng human-related features, in line with the findings of Wen and Zhou (2024). Images and POIs offer essential additional informat8ion, boosting performance across both types of tasks. Overall, MoRA achieves the best performance in every domain and downstream task, validating the effectiveness of its multimodal fusion approach.
Table 6: Ablation on auxiliary modalities in mobility-as-backbone fusion.
Social Tasks Economic Tasks AVG $R^{2}$ $\uparrow$
POP EDU ELD HSR CRI NTL HOU ENE COS
w/o POI 0.835 0.835 0.796 0.800 0.755 0.590 0.702 0.680 0.895 0.765
w/o Image 0.832 0.841 0.791 0.803 0.755 0.599 0.673 0.665 0.900 0.762
w/o Demo 0.824 0.827 0.765 0.793 0.753 0.605 0.685 0.590 0.893 0.748
MoRA 0.829 0.853 0.812 0.810 0.759 0.615 0.703 0.719 0.909 0.779
A.6 Sensitivity to H3 Grid Resolution
We extend our study at H3 resolution level 6 to its adjacent levels, 5 and 7, to evaluate the model’s performance across different spatial scales. Compared to level 6 H3 cells, which cover an average area of 36.13 km², cells at levels 5 and 7 cover approximately 253.61 km² and 5.18 km², respectively. The East China region is selected for testing to facilitate ease of operation without losing generalizability. We reconstruct the training dataset for levels 5 and 7 from scratch, including forming trajectory-based mobility graphs, gathering satellite imagery and POIs for encodings, and processing demographic distributions data. Three separate experiments are conducted, with corresponding sets of downstream data processed for evaluation. Since cities in East China are limited in quantity, we choose Ridge Regression for energy consumption prediction across all levels to ensure consistency. For the remaining tasks, LightGBM is adopted. Table 7 reports model performance with varying spatial resolutions. Overall, performance generally improves as spatial granularity increases, though the rationale varies by the level of the downstream tasks. For grid-level predictions, finer resolution can enhance localized prediction. However, an exception is observed with crime prediction, which follows a reversed pattern. This might be explained by the fact that crime becomes more heterogeneous and exhibits greater variations at finer scales. At the county and city levels, where aggregation smooths out local noise and highlights more significant features, performance improves with grid granularity.
Table 7: Sensitivity of model performance using different grid resolutions. (R²)
@@ A.6 Sensitivity to H3 Grid Resolution
A.7 The Use of Large Language Models (LLMs)
We employed a large language model (LLM) exclusively for language polishing, including grammar correction, stylistic refinement, and improvements to clarity and readability. The LLM did not contribute to the conception of the study, data collection, model design, statistical analysis, or interpretation of results, nor did it generate figures, tables, or substantive content. All methodological choices, analyses, and conclusions were conceived, executed, and validated by the authors. Drafts edited with the LLM were subsequently reviewed line by line by the authors to ensure technical accuracy, fidelity to the original meaning, and alignment with ethical and scholarly standards. The authors bear full responsibility for all scientific content presented in this work.